evaluale the iterated internal g 1 2 V4-22 - dy dz dy X + D
Answer #1

d 41 dydzdu m. jam dzak 2 de du 2 254-zz azdu И +1 how, net, 4-2²=+² - t = 14-22 -22 d2 =2tdt Zd2 = -tdt = 1 p + (-tdt) du o

o Itn = 3 [In(ital] = 3 (1n2-1n1) = 3 1n2

please upvote...thank u:)